Please help me solve: How much fuel that sells for $.85 per gallon should be added to 13 gallons of fuel that sells for $1.75 per gallon to obtain a mixture that sells for $1.50 per gallon?
Thanks

Answer by lwsshak3(11628)

Please help me solve: How much fuel that sells for $.85 per gallon should be added to 13 gallons of fuel that sells for $1.75 per gallon to obtain a mixture that sells for $1.50 per gallon?
**
let x=gallons of $.85 fuel to be added.
13+x=gallons of mixture
.85x+1.75*13=1.50(13+x)
.85x+22.75=19.50+1.5x
.65x=3.25
x=3.25/.65=5
gallons of $.85 fuel to be added=5